### v2

# VERSA: A Versatile Evaluation Toolkit for Speech, Audio, and Music

Abstract
In this work, we introduce VERSA, a unified and standardized evaluation toolkit designed for various speech, audio, and music signals. The toolkit features a Pythonic interface with flexible configuration and dependency control, making it user-friendly and efficient. With full installation, VERSA offers 65 metrics with 729 metric variations based on different configurations. These metrics encompass evaluations utilizing diverse external resources, including matching and non-matching reference audio, text transcriptions, and text captions. As a lightweight yet comprehensive toolkit, VERSA is versatile to support the evaluation of a wide range of downstream scenarios. To demonstrate its capabilities, this work highlights example use cases for VERSA, including audio coding, speech synthesis, speech enhancement, singing synthesis, and music generation. The toolkit is available at https://github.com/wavlab-speech/versa.
VERSA: A Versatile Evaluation Toolkit for Speech, Audio, and Music
Jiatong Shi1, Hyejin Shim1, Jinchuan Tian1, Siddhant Arora1, Haibin Wu2, Darius Petermann3, Jia Qi Yip4, You Zhang5, Yuxun Tang6, Wangyou Zhang7, Dareen Alharthi1, Yichen Huang1, Koichi Saito8, Jionghao Han1, Yiwen Zhao1, Chris Donahue1, Shinji Watanabe1, 1Carnegie Mellon University, 2Microsoft, 3Indiana University, 4Nanyang Technological University, 5University of Rochester, 6Renmin University of China, 7Shanghai Jiaotong University, 8Sony AI
1 Introduction
With the rapid advancements in artificial intelligence-generated content (AIGC), deep generative models have demonstrated remarkable capabilities in producing high-quality outputs across various domains, including image, video, and sound generation. As generative models become increasingly sophisticated, the need for comprehensive AIGC evaluation has grown, aimed at identifying the strengths and weaknesses of the generated outputs.
As an essential part of the language processing community, diverse generative models for speech, music, and general audio have shown significant potential in applications such as conversational interfaces McTear (2002), entertainment Fraser et al. (2018); Fancourt and Steptoe (2019); Dash and Agres (2024), and task management Kulkarni et al. (2019). Due to the perceptual nature of sound-based applications, human subjective assessment is widely regarded as the gold standard for evaluating sound generative models.
The most fundamental and widely used metric for these models is the mean opinion score (MOS) Recommendation (1994). The initial purpose of MOS was to measure the naturalness of generated audio, but it has since evolved into various specific forms, such as evaluating speaker similarity Toda et al. (2016), comparative performance against baseline systems Harada and Ohmuro (1999), emotional similarity Choi et al. (2019), and alignment with prompts Guo et al. (2023); Li et al. (2024). Despite its importance, subjective evaluations relying on human input are challenging to conduct due to their labor-intensive nature. Furthermore, achieving statistically significant results often requires a substantial number of samples Rosenberg and Ramabhadran (2017). Additionally, range-equalizing bias is frequently observed in MOS evaluations due to the psychological grounding of human subjective assessments Cooper and Yamagishi (2023); Le Maguer et al. (2024). Such biases introduce considerable challenges in achieving comparable results across different evaluation datasets and participants, thereby complicating the process of benchmarking generative models effectively. Last but not least, certain evaluation variants may require individuals with expert knowledge to conduct the assessment, particularly in the music domain Ji et al. (2020).
An alternative approach is to develop automatic metrics that align with human preferences. The design of such metrics can vary based on the use of external references and the specific application scenarios, which further complicates their selection.
The basic setup for evaluation involves using only the candidate audio being evaluated Liang and Kubichek (1994); Falk and Chan (2006); Lo et al. (2019); Saeki et al. (2022), which has been actively discussed in recent VoiceMOS challenges Huang et al. (2022); Cooper et al. (2023); Huang et al. (2024b). In addition, a variety of external resources can be optionally utilized for evaluation, as illustrated in Fig.1. These resources may include matching reference audio Rix et al. (2001), non-matching reference audio Manocha et al. (2021); Ragano et al. (2024a), textual transcription Hayashi et al. (2020), textual captions Huang et al. (2023), or visual cues Hu et al. (2021). [^1] Note that we specifically exclude pre-trained models from our notion of “external resources”.
Figure 1: Using various external resources for automatic sound evaluation. External resources include any matching reference signals, non-matching reference signals, transcriptions, visual cues, or textual captions.
The focus of automatic metrics can vary significantly depending on the application scenario. For instance, in voice conversion, discriminative speaker embeddings can be employed to measure speaker similarity between the generated speech and speech from the same speakers Das et al. (2020). In cases involving domain-specific content, such as singing, pre-trained models in singing voices may better align with assessments of singing naturalness Tang et al. (2024). For a fine-grained, sample-level generation, signal-to-noise ratio-related metrics are more suitable, particularly for tasks such as speech enhancement and separation Luo and Mesgarani (2019). Due to the creative nature of music, distributional metrics are often better suited for evaluating music generation Kilgour et al. (2019). Given the diversity of audio signals, effectively evaluating them requires a broad understanding of general audio events and their contextual relevance Huang et al. (2023).
Considering the diversity of metrics and the evolution of generative models for speech, audio, and music, the need for standardized evaluation metrics has become increasingly evident. Without a unified framework, the diversity in evaluation methodologies often leads to inconsistent results, making it challenging to benchmark models or assess advancements effectively. Standardization ensures that all systems are evaluated under comparable conditions, fostering fairness, reproducibility, and meaningful insights across studies. Moreover, centralizing metrics within a single toolkit not only reduces redundancy and inefficiencies but also encourages collaboration by providing researchers with a shared foundation for assessing performance. This need for consistency and centralization underpins the development of VERSA, a toolkit designed to address these challenges.
Extended from its prior version in Shi et al. (2024), this work introduces a versatile evaluation toolkit for speech, audio, and music: VERSA. Built on a Pythonic interface, VERSA integrates 65 metrics and more than 729 variants, offering a wide array of automatic evaluation tools tailored for speech, audio, and music. By providing diverse metrics, VERSA aims to serve as a one-stop solution for multi-domain, multi-level, and multi-focus sound evaluation across various downstream applications. With examples showcasing the use of VERSA in ESPnet-Codec and ESPnet-SpeechLM, we anticipate that the VERSA toolkit will become a key component in advancing sound generation benchmarks, addressing challenges in speech generation, and supporting multi-modal generative frameworks. The codebase is publicly available at https://github.com/wavlab-speech/versa. [^2] A video demonstration is available at https://youtu.be/t7UP1uFvaCM.
2 VERSA
This section details the design of VERSA, focusing on its general framework, supported metrics, and potential benefits for the community.
2.1 VERSA Framework
Figure 2: Directory structure of VERSA. Detailed discussion can be found in Sec. 2.1
As illustrated in Fig. 2, the core library of VERSA is implemented as a Python package with two straightforward interfaces: scorer.py and aggregate_result.py. The scorer.py interface computes the automatic metrics, while aggregate_result.py consolidates the results into a final report for users.
Once installed via pip, using VERSA is as simple as the following:
1python versa/bin/scorer.py \
2 --score_config egs/general.yaml \
3 --gt <ground truth audio list> \
4 --pred <candidate audio list> \
5 --output_file test_result
Listing 1: A simple scorer interface.
where the ground truth audio list is optional and is not required for independent metrics.
I/O Interface: VERSA offers three I/O interfaces for handling audio samples: the Soundfile interface, the Directory interface, and the Kaldi interface. These interfaces support a variety of audio formats (e.g., PCM, FLAC, MP3, Kaldi-ARK) and different file organizations, such as wav.scp files or individual audio files stored within a parent directory. For each (candidate, reference) audio signal pair, a resampling is conducted for each metric according to the specific sampling rate required by that metric. librosa is used for resampling.
Flexible Configuration: VERSA employs a unified YAML-style configuration file to define which metrics to use and control their detailed setups. While users can directly explore the library code, we also provide example configuration files for different metrics in the egs directory, as shown in Fig. 2. For instance, egs/general.yaml provides a configuration template for the default installation metrics. Additionally, individual YAML configuration files for specific metrics are available under egs/separate_metric. Some example configurations are discussed in Appendix A.
Strict Dependency Control: Managing dependencies can be challenging when using diverse evaluation metrics. To address this, VERSA offers a minimal-dependency installation that supports a core set of metrics by default, while additional installation scripts are provided for metrics with extra requirements. This approach significantly reduces dependency overhead during VERSA installation, especially for metrics with heavy dependencies or complex compilation needs. As shown in Fig. 2, these additional installation scripts are located in the tools directory.
To ensure correct model usage, many official packages released by model providers enforce strict dependency control by specifying exact package versions (e.g., specific versions of PyTorch or NumPy). While this ensures compatibility, it often introduces unnecessary dependency conflicts with major packages. To provide a more flexible environment for users, VERSA bypasses these strict version requirements by adapting the interfaces of such metrics into our own local forks. These forks are supplemented with additional numerical tests to ensure functionality without adhering to rigid version control.
Moreover, using our own fork allows us to integrate VERSA-specific interfaces into external metrics that may otherwise conflict with the toolkit’s design philosophy. This flexibility ensures a consistent and seamless interface across the VERSA library, enhancing usability and maintaining the toolkit’s design integrity.
In Appendix B, we further discuss additional system design concepts on job scheduling, test protocol, resource management, and community-driven contribution guidelines.
Table 1: List of supported metrics in VERSA. The “Base" column indicates whether the metrics are included in the minimum installation of VERSA. The “Model Based” column represents metrics that need pre-trained models. The “Target Direction” column indicates which direction is desirable for each metric without being overly technical.
No. Type Domain Name Base Variants Range Model Based Target Direction Reference
Speech Audio Music
1 Independent Deep Noise Suppression MOS Score of P.835 (DNSMOS P.835) 1 [1, 5] $\uparrow$ Reddy et al. (2022)
2 Deep Noise Suppression MOS Score of P.808 (DNSMOS P.808) 1 [1, 5] $\uparrow$ Reddy et al. (2021)
3 Speech Quality and Naturalness Assessment (NISQA) 3 [1, 5] $\uparrow$ Mittag et al. (2021)
4 UTokyo-SaruLab System for VoiceMOS 2022 (UTMOS) 1 [1, 5] $\uparrow$ Saeki et al. (2022)
5 Packet Loss Concealment-focus MOS (PLCMOS) 1 [1, 5] $\uparrow$ Diener et al. (2023)
6 Torch-Squim PESQ (TS-PESQ) 1 [1, 5] $\uparrow$ Kumar et al. (2023)
7 Torch-Squim STOI (TS-STOI) 1 [0, 1] $\uparrow$ Kumar et al. (2023)
8 Torch-Squim SI-SNR (TS-SI-SNR) 1 (-inf, inf) $\uparrow$ Kumar et al. (2023)
9 Singing voice MOS (SingMOS) 1 [1, 5] $\uparrow$ Tang et al. (2024)
10 Subjective Speech Quality Assessment (SSQA) in SHEET Toolkit 1 [1, 5] $\uparrow$ Huang et al. (2024a)
11 UTokyo-SaruLab System for VoiceMOS 2024 (UTMOSv2) 1 [1, 5] $\uparrow$ Baba et al. (2024)
12 Speech Quality with Contrastive Regression (SCOREQ) wo. Ref. 2 [1, 5] $\uparrow$ Ragano et al. (2024b)
13 Speech Enhancement-based SI-SNR (SE-SI-SNR) 69 (-inf, inf) $\uparrow$ Zhang et al. (2024a)
14 Speech Enhancement-based CI-SDR (SE-CI-SDR) 69 (-inf, inf) $\uparrow$ Zhang et al. (2024a)
15 Speech Enhancement-based SAR (SE-SAR) 69 (-inf, inf) $\uparrow$ Zhang et al. (2024a)
16 Speech Enhancement-based SDR (SE-SDR) 69 (-inf, inf) $\uparrow$ Zhang et al. (2024a)
17 Prompting Audio-Language Models (PAM) metric 1 [0, 1] $\uparrow$ Deshmukh et al. (2024)
18 Speech-to-reverberation Modulation Energy Ratio (SRMR) 1 [0, inf) $\uparrow$ Falk et al. (2010)
19 Voice Activity Detection (VAD) 1 - - SileroTeam (2024)
20 Speaker Turn Taking (SPK-TT) 1 [0, inf) - Goodwin and Heritage (1990)
21 Speaking Word Rate (SWR) 7 [0, inf) - Radford et al. (2023)
22 Anti-spoofing Score (SpoofS) 2 [0, 1] $\uparrow$ Jung et al. (2021)
23 Language Identification (LID) 17 - - Peng et al. (2023)
24 Audiobox Aesthetics (AA) 1 [1, 10] $\uparrow$ Tjandra et al. (2025)
25 Dependent Mel Cepstral Distortion (MCD) 1 [0, inf) $\downarrow$ Kubichek (1993)
26 F0 Correlation (F0-CORR) 1 [-1, 1] $\uparrow$ Hayashi et al. (2021)
27 F0 Root Mean Square Error (F0-RMSE) 1 [0, inf) $\downarrow$ Hayashi et al. (2021)
28 Signal-to-interference Ratio (SIR) 1 (-inf, inf) $\uparrow$ Févotte et al. (2005)
29 Signal-to-artifact Ratio (SAR) 1 (-inf, inf) $\uparrow$ Févotte et al. (2005)
30 Signal-to-distortion Ratio (SDR) 1 (-inf, inf) $\uparrow$ Févotte et al. (2005)
31 Convolutional scale-invariant signal-to-distortion ratio (CI-SDR) 1 (-inf, inf) $\uparrow$ Boeddeker et al. (2021)
32 Scale-invariant signal-to-noise ratio (SI-SNR) 1 (-inf, inf) $\uparrow$ Luo and Mesgarani (2018)
33 Perceptual Evaluation of Speech Quality (PESQ) 1 [1, 5] $\uparrow$ Rix et al. (2001)
34 Short-Time Objective Intelligibility (STOI) 1 [0, 1] $\uparrow$ Taal et al. (2011)
35 Speech BERT Score (D-BERT) 1 [-1, 1] $\uparrow$ Saeki et al. (2024)
36 Discrete Speech BLEU Score (D-BLEU) 1 [0, 1] $\uparrow$ Saeki et al. (2024)
37 Discrete Speech Token Edit Distance (D-Distance) 1 [0, 1] $\uparrow$ Saeki et al. (2024)
38 Dynamic Time Warping Cost (WARP-Q) 1 [0, inf) $\uparrow$ Jassim et al. (2022)
39 Speech Quality with Contrastive Regression (SCOREQ) w. Ref. 2 [1, 5] $\uparrow$ Ragano et al. (2024b)
40 2f-Model 1 [0, 100] $\uparrow$ Kastner and Herre (2019)
41 Log-weighted Mean Square Error (Log-WMSE) 1 (-inf, inf) $\uparrow$ Jordal (2023)
42 ASR-oriented Mismatch Error Rate (ASR-Mismatch) 7 [0, inf] $\downarrow$ Radford et al. (2023)
43 Virtual Speech Quality Objective Listener (VISQOL) 4 [1,5] $\uparrow$ Chinen et al. (2020)
44 Frequency-Weighted SEGmental SNR (FWSEGSNR) 1 (-inf, inf) $\uparrow$ Tribolet et al. (1978)
45 Weighted Spectral Slope (WSS) 1 [0, inf) $\downarrow$ Klatt (1982)
46 Cepstrum Distance (CD) 1 [0, inf) $\downarrow$ Barnwell III et al. (1988)
47 Composite Objective Speech Quality (Csig, Cbak, Covl) 1 [1, 5] $\uparrow$ Hu and Loizou (2007)
48 Coherence and Speech Intelligibility Index (CSII) 1 [0, 1] $\uparrow$ Kates and Arehart (2005)
49 Normalized-Covariance Measure (NCM) 1 [-1, 1] $\uparrow$ Chen and Loizou (2010)
50 Non-match Non-matching Reference Speech Quality Assessment (Noresqa) 2 [1, 5] $\uparrow$ Manocha et al. (2021)
51 Torch-Squim MOS (TS-MOS) 1 [1, 5] $\uparrow$ Kumar et al. (2023)
52 ESPnet ASR Model Word Error Rate (ESPnet-WER) 270 [0, inf) $\downarrow$ Watanabe et al. (2018)
53 ESPnet OWSM Model Word Error Rate (OWSM-WER) 17 [0, inf) $\downarrow$ Peng et al. (2023)
54 OpenAI Whisper Model Word Error Rate (Whisper-WER) 7 [0, inf) $\downarrow$ Radford et al. (2023)
55 Emotion Similarity (EMO-SIM) 1 [-1, 1] $\uparrow$ Ma et al. (2024)
56 Speaker Similarity (SPK-SIM) 15 [-1, 1] $\uparrow$ weon Jung et al. (2024)
57 Non-Matching Reference Audio Quality Assessment (NOMAD) 1 [1, 5] $\uparrow$ Ragano et al. (2024a)
58 Contrastive Language-Audio Pretraining Score (CLAP Score) 1 [-1, 1] $\uparrow$ Huang et al. (2023)
59 Accompaniment Prompt Adherence (APA) 1 [-1, 1] $\uparrow$ Grachten (2024)
60 Log Likelihood Ratio (LLR) 1 [0, inf) $\uparrow$ Hu and Loizou (2007)
61 Distributional Fréchet Audio Distance Audio Distance (FAD) 11 [0, inf) $\downarrow$ Roblek et al. (2019)
62 Kullback-Leibler Divergence on Embedding Distribution (KLD) 11 [0, inf) $\downarrow$ Roblek et al. (2019)
63 Density in Embedding Space (Density-Embedding) 11 [0, inf) $\uparrow$ Naeem et al. (2020)
64 Coverage in Embedding Space (Coverage-Embedding) 11 [0, 1] $\uparrow$ Naeem et al. (2020)
65 Kernel Distance/Maximum Mean Discrepancy (KID) 11 [0, inf) $\downarrow$ Roblek et al. (2019)
- Total 56 23 23 - 40 729 - 48 -
Table 2: Comparison to related toolkits. The number of metrics are collected at 12/09/2024.
Name Domain Open-source Link Metric Type
Speech Audio Music
ESPnet Watanabe et al. (2018) https://github.com/espnet/espnet 16
Apmhion Zhang et al. (2023a) https://github.com/open-mmlab/Amphion 15
SHEET Huang et al. (2024a) https://github.com/unilight/sheet 3
SpeechMOS https://pypi.org/project/speechmos 3
BSS-EVAL Févotte et al. (2005) https://pypi.org/project/fast-bss-eval 4
ClearerVoice-SpeechScore https://github.com/modelscope/ClearerVoice-Studio 14
AudioLDM-Eval Liu et al. (2023) https://github.com/haoheliu/audioldm_eval 9
Stable-Audio-Metric https://github.com/Stability-AI/stable-audio-metrics 3
Sony Audio-Metrics Grachten (2024) https://github.com/SonyCSLParis/audio-metrics 4
FADTK Gui et al. (2024) https://github.com/microsoft/fadtk 11
Pysepm https://github.com/schmiph2/pysepm 10
AudioCraft Copet et al. (2024) https://github.com/facebookresearch/audiocraft/blob/main/docs/METRICS.md 6
AutaTK Vinay and Lerch (2023) https://github.com/Ashvala/AQUA-Tk 9
VERSA https://github.com/shinjiwlab/versa 65
2.2 Supported Metrics
VERSA stands out for its extensive range of supported metrics, categorized into four main types:
Independent metrics: These metrics do not require any dependent external resources, other than pre-trained models. Notably, we adopt an extended form of independent metrics in speech and audio assessment, which also considers profiling-related metrics, such as voice activity detection and speaker turn-taking. VERSA currently supports 22 independent metrics.
Dependent metrics: These metrics rely on matching sound references. In VERSA, 25 dependent metrics are supported.
Non-matching metrics: These metrics use non-matching reference data or different modalities. VERSA currently supports 11 non-matching metrics.
Distributional metrics: These metrics conduct distribution comparisons between two datasets, providing a more general view of the generative models’ performance. VERSA currently supports 5 distributional metrics.
As summarized in Table 1, VERSA supports a total of 65 metrics, 39 of which are included in the minimal installation. Of these, 54 metrics are applicable to speech tasks, 22 to general audio tasks, and 22 to music tasks. Additionally, several metrics feature variations based on different pre-trained models, such as word error rate, speaker similarity, and Fréchet Audio Distance (FAD) scores. By simply modifying the configuration file, [^3] An example of a configuration change is provided in Appendix A.2. VERSA can generate up to 729 distinct metric variants, offering flexibility for a wide range of evaluation scenarios.
2.3 Advantages of VERSA
This section highlights the key benefits of VERSA, focusing on its ability to ensure consistency, facilitate comparability, provide comprehensive insights, and enhance efficiency.
Consistency: VERSA ensures uniform evaluation criteria across experiments, addressing a critical need in the field of sound generative models. By standardizing the implementation of metrics, VERSA minimizes the variability introduced by subjective judgments or inconsistent evaluation setups (e.g., coding environments).
Comparability: One of VERSA’s key advantages is its ability to facilitate benchmarking against existing models and methodologies. By providing a unified set of metrics, it ensures that evaluations are conducted under fair and objective conditions. This comparability is important for assessing the relative performance of new approaches, fostering innovation, and enabling the broader research community to progress effectively.
Comprehensiveness: VERSA supports a wide array of evaluation metrics, including dimensions such as perceptual quality, intelligibility, affective information, and creative diversity. By incorporating these diverse measures, the toolkit provides a holistic view of system performance, especially for researchers to gain deeper insights into both the strengths and weaknesses of each method.
Efficiency: With its all-in-one design, VERSA significantly enhances efficiency by supporting multiple metrics within a single toolkit. Users no longer need to rely on separate tools or perform manual calculations to assess different aspects of audio performance. This workflow reduces the time, effort, and potential errors associated with using fragmented evaluation methods, accelerating the overall research and development process.
3 Comparison to Other Toolkits
As discussed in Sec. 1, the challenges associated with subjective evaluation have propelled the community toward exploring objective evaluation metrics. The growing demand for sound evaluation toolkits has led to numerous efforts in domain-specific evaluation for sound generation.
In the speech domain, prior to the deep learning era, the International Telecommunication Union Telecommunication Standardization Sector (ITU-T) played a pivotal role in designing evaluation metrics for speech processing tasks such as speech coding and speech enhancement. More recently, text-to-speech (TTS) toolkits have begun incorporating speech quality assessment features, exemplified by ESPnet-TTS Hayashi et al. (2020, 2021) and Amphion Zhang et al. (2023a). Additionally, the Speech Human Evaluation Estimation Toolkit (SHEET) framework provides an all-in-one recipe-style toolkit for data preparation, speech quality prediction model training, and evaluation Huang et al. (2024a). In speech enhancement, foundational signal-level metrics were first consolidated in Févotte et al. (2005), followed by extensions such as Pysepm, which supports 10 metrics Loizou (2013). More recent advancements include the addition of 14 speech enhancement metrics in ClearerVoice ClearerVoice-Studio (2023).
In the audio domain, AudioLDM-Eval focuses on evaluating audio language models with nine types of metrics Liu et al. (2023, 2024). Stability AI has introduced three audio metrics Stability-AI (2023), while Sony CSL has open-sourced four additional types of audio metrics Grachten (2024).
In the music domain, MIR_EVAL is a pioneering toolkit that aggregates metrics for music information retrieval tasks Raffel et al. (2014). More recently, Microsoft released FADTK, which emphasizes a comprehensive FAD-embedding space for generative music evaluation Gui et al. (2024).
A summary of the related toolkits is presented in Table 2. Each framework has made significant contributions to the community, many serving as foundational tools for sound generation research. Building on these previous toolkits, VERSA distinguishes itself with a general design applicable across multiple domains and its comprehensive inclusion of 65 metrics with 729 variants—capabilities that have not been achieved before.
4 Demonstration
We demonstrate several use cases of VERSA in diverse scenarios, including speech coding, speech synthesis, speech enhancement, singing synthesis, and music generation.
Speech coding remains one of the most widely utilized applications within the speech-processing community. In this demonstration, we leverage VERSA to evaluate nine publicly available codecs. More details and corresponding results are discussed in Table 4 in Appendix C.1.
Text-to-speech aims to convert text into speech signals. In this demonstration, we use VERSA to compare nine open-source TTS models. More details and corresponding are shown in Table 6, Appendix C.2.
Speech enhancement targets the improvement of speech transmission in different environments. In this paper, we also demonstrate the VERSA usage in speech enhancement in three speech enhancement models. More details and corresponding information are shown in Table 7, Appendix C.3.
Singing synthesis is an intersection between speech and music generation, where both speech-oriented and music-oriented evaluation metrics are needed. VERSA offers a one-stop solution to this problem based on the collection of a variety of metrics in each domain. In this work, we demonstrate the evaluation of a range of singing voice synthesis models. Table 9, Appendix C.4 shows more details and corresponding information.
Music generation and its evaluation have received increasing attention due to the rapid progress in model development. To consider both creativity and musical harmony, recent evaluation methods mostly utilize distribution metrics for the evaluation, exhibiting a large difference to other sound generation systems. In this demonstration, VERSA aggregates a range of music generation evaluation metrics into a single toolkit. More details and corresponding information are shown in Table 11, Appendix C.5.
5 Conclusion
In this work, we introduced VERSA, a comprehensive and versatile evaluation toolkit for assessing speech, audio, and music signals. With its flexible Pythonic interface and extensive suite of metrics, VERSA empowers researchers and developers to conduct rigorous and reproducible evaluations across diverse generative tasks.
Through its integration of more than 65 metrics and 729 variants, VERSA provides unparalleled support for evaluating speech synthesis, audio coding, music generation, and more. The toolkit not only simplifies the evaluation process but also addresses challenges such as bias in subjective evaluations and the need for domain-specific expertise.
6 Ethics Statement
VERSA is designed to address the challenges of evaluating sound generative models in diverse linguistic, cultural, and acoustic contexts. Generative sound models often risk perpetuating biases due to their reliance on datasets that may underrepresent certain languages, accents, and cultural expressions. To mitigate these risks, VERSA incorporates evaluation metrics and methodologies that accommodate a wide range of audio characteristics, including tonal, phonetic, and rhythmic variations present across global languages and music traditions.
The toolkit encourages the use of regionally diverse datasets and provides flexibility to integrate culturally specific evaluation resources, such as non-standard phonetic systems or traditional musical structures. By doing so, VERSA seeks to foster the development of sound generative models that respect and represent the full spectrum of human audio diversity.
Through these efforts, VERSA empowers researchers and developers to create more inclusive generative models, ensuring that advances in speech, audio, and music technologies benefit communities worldwide, regardless of linguistic or cultural background.
7 Broader Impact
7.1 Current Limitations
While VERSA offers a comprehensive and versatile evaluation framework for speech, audio, and music generation, it is not without its limitations. Below, we outline some areas where the toolkit could be further improved:
Dependence on External Resources: Many of VERSA’s metrics require external resources, such as pre-trained models, reference datasets, or additional Python packages. The quality and diversity of these resources can impact the accuracy and fairness of evaluations, particularly for underrepresented languages and cultures. While VERSA provides flexible configurations to accommodate various scenarios, the availability of such resources remains a bottleneck in some cases.
Bias in Metric Design: Despite efforts to include diverse evaluation metrics, some metrics may still reflect biases inherent in the training data or methodologies used to develop them. For example, evaluation frameworks optimized for widely spoken languages or Western music may not fully capture the nuances of less-studied languages, dialects, or musical traditions. This bias can lead to less accurate evaluations for certain domains or cultural contexts.
Subjectivity in Perceptual Metrics: While VERSA incorporates automatic metrics designed to align with human subjective assessments, these metrics may not always perfectly reflect human preferences or perceptions. Human evaluations remain the gold standard for certain aspects of sound quality, naturalness, and emotional expressiveness, which automated metrics cannot fully replicate.
Evolving Standards in Generative Models: The rapid advancement of generative audio technologies means that new evaluation needs and metrics may arise that VERSA does not yet support. As a result, maintaining the toolkit’s relevance and adaptability will require ongoing updates and contributions from the community.
7.2 Future Adaptation
By accommodating a wide range of configurations, external resources, and application scenarios, we expect VERSA to bridge the gap between human subjective assessments and automatic evaluation metrics, ensuring robust and scalable benchmarking. The adoption and usage of VERSA are not restricted by geographic boundaries. As an open-source toolkit, it is accessible globally to researchers and practitioners, fostering international collaboration in advancing sound generation technologies. VERSA is designed with adaptability in mind, allowing users to integrate their local resources and datasets to perform culturally and regionally relevant evaluations.
Furthermore, VERSA facilitates cross-border innovation by providing a standardized framework for evaluating generative audio models. This standardization reduces duplication of effort and promotes reproducibility, ensuring that advancements in sound generation and evaluation can transcend political and cultural borders. Our commitment to accessibility and inclusivity reinforces our belief in the universal potential of AI to benefit humanity without being limited by geographic, linguistic, or cultural barriers.
Looking ahead, we envision VERSA as a key enabler for advancing the field of sound generation. Its modular design ensures adaptability to future developments, while its open-source availability fosters collaboration and community-driven enhancements. By setting a new standard for sound evaluation, VERSA could pave the way for more transparent and effective comparisons of generative models, ultimately accelerating progress in AI-driven audio and music technologies.
Acknowledgements
This work is supported by the Defence Science and Technology Agency (DSTA) in Singapore. We would like to thank Daniel Leong and Megan Choo for their valuable comments. Experiments of this work used the Bridges2 at PSC and Delta/DeltaAI NCSA computing systems through allocation CIS210014 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, supported by National Science Foundation grants 2138259, 2138286, 2138307, 2137603, and 2138296.
Appendix A Example Configuration
A.1 Base Configuration
We demonstrate an example configuration file in Listing 2, illustrating the ease of setup and customization within VERSA. The configuration options are designed to cater to various user needs, offering flexibility for advanced users to tailor evaluations to specific scenarios.
Simultaneously, default settings are provided to ensure an intuitive experience for new users, allowing them to quickly begin leveraging the toolkit without requiring extensive configuration knowledge. These thoughtful defaults strike a balance between simplicity and functionality, empowering both novice and experienced users to perform robust evaluations effortlessly.
# Example YAML config
# mcd f0 related metrics
# -- mcd: mel cepstral distortion
# -- f0_corr: f0 correlation
# -- f0_rmse: f0 root mean square error
- name: mcd_f0
f0min: 40
f0max: 800
mcep_shift: 5
mcep_fftl: 1024
mcep_dim: 39
mcep_alpha: 0.466
seq_mismatch_tolerance: 0.1
power_threshold: -20
dtw: false
# signal related metrics
# -- sir: signal to interference ratio
# -- sar: signal to artifact ratio
# -- sdr: signal to distortion ratio
# -- ci-sdr: scale-invariant signal to distortion ratio
# -- si-snri: scale-invariant signal to noise ratio improvement
- name: signal_metric
# pesq related metrics
# -- pesq: perceptual evaluation of speech quality
- name: pesq
# stoi related metrics
# -- stoi: short-time objective intelligibility
- name: stoi
# discrete speech metrics
# -- speech_bert: speech bert score
# -- speech_bleu: speech bleu score
# -- speech_token_distance: speech token distance score
- name: discrete_speech
# pseudo subjective metrics
# -- utmos: UT-MOS score
# -- dnsmos: DNS-MOS score
# -- plcmos: PLC-MOS score
- name: pseudo_mos
predictor_types: ["utmos", "dnsmos", "plcmos", "singmos"]
predictor_args:
utmos:
fs: 16000
dnsmos:
fs: 16000
plcmos:
fs: 16000
# speaker related metrics
# -- spk_similarity: speaker cosine similarity
# model tag can be any ESPnet-SPK huggingface repo at
# https://huggingface.co/espnet
- name: speaker
model_tag: default
# torchaudio-squim
# -- torch_squim_pesq: reference-less pesq
# -- torch_squim_stoi: reference-less stoi
# -- torch_squim_si_sdr: reference-less si-sdr
# -- torch_squim_mos: MOS score with reference
- name: squim_ref
- name: squim_no_ref
# An overall model on MOS-bench from Sheet toolkit
# More info in https://github.com/unilight/sheet/tree/main
# --sheet_ssqa: the mos prediction from sheet_ssqa
- name: sheet_ssqa
# Word error rate with OpenAI-Whisper model
# More model_tag can be from https://github.com/openai/whisper?tab=readme-ov-file#available-models-and-languages .
# The default model is ‘large-v3‘.
# NOTE(jiatong): further aggregation are necessary for corpus-level WER/CER
# --whisper_hyp_text: the hypothesis from ESPnet ASR decoding
# --ref_text: reference text (after cleaner)
# --whisper_wer_delete: delete errors
# --whisper_wer_insert: insertion errors
# --whisper_wer_replace: replacement errors
# --whisper_wer_equal: correct matching words/character counts
# --whisper_cer_delete: delete errors
# --whisper_cer_insert: insertion errors
# --whisper_cer_replace: replacement errors
# --whisper_cer_equal: correct matching words/character counts
- name: whisper_wer
model_tag: default
beam_size: 5
text_cleaner: whisper_basic
# Speech Enhancement-based Metrics
# model tag can be any ESPnet-SE huggingface repo
# -- se_sdr: the SDR from a reference speech enhancement model
# -- se_sar: the SAR from a reference speech enhancement model
# -- se_ci_sdr: the CI-SDR from a reference speech enhancement model
# -- se_si_snr: the SI-SNR from a rerference speech enhancement model
- name: se_snr
model_tag: default
Listing 2: An example of the configuration file.
A.2 Configuration Change for Model Variations
As mentioned in Sec.2.2, the configuration can be easily adjusted to accommodate different model variations. For instance, by default, we use the Rawnet-based speaker embedding weon Jung et al. (2024) for speaker similarity calculation, as demonstrated in Listing 3.
1# Example Speaker Similarity YAML config
2
3# speaker related metrics
4# -- spk_similarity: speaker cosine similarity
5# model tag can be any ESPnet-SPK huggingface repo at
6# https://huggingface.co/espnet
7- name: speaker
8 model_tag: default
Listing 3: An example of the speaker similarity configuration file.
To switch the backend speaker embedding model, you can simply modify the model_tag to any supported speaker embedding model available in the ESPnet Huggingface collection. [^4] The complete list can be found at https://huggingface.co/models?other=speaker-recognition&sort=trending&search=espnet. An example is provided in Listing 4.
1# Example Speaker Similarity YAML config
2
3# speaker related metrics
4# -- spk_similarity: speaker cosine similarity
5- name: speaker
6 model_tag: espnet/voxcelebs12_ecapa_wavlm_joint
Listing 4: An example of the speaker similarity configuration file with a switched base model.
Appendix B Additional System Design
Job-scheduling Systems: In VERSA’s main library, Python multi-processing is not natively supported to avoid complicating the codebase design. Instead, VERSA achieves multi-processing through job-scheduling systems. Inspired by the approaches used in Kaldi and ESPnet Povey et al. (2011); Watanabe et al. (2018), this design enables efficient task scheduling on both computing clusters and local operating systems. To facilitate this, VERSA includes an example script, launch.sh, which utilizes SLURM as the job-scheduling system to manage multi-processing tasks effectively.
Test Protocol with Continuous Integration: To ensure toolkit consistency and guarantee numerical stability across diverse scenarios, VERSA is actively integrating to continuous integration testing. Unit tests are implemented for each metric, supporting a range of Python versions and external package dependencies.
Table 3: Detailed information of codec models and inference setups.
Codec kBPS Num. Levels Token Rate Frame Rate Sampling Rate Link
Encodec Défossez et al. (2024a) 6 8 600 75 24kHz (https://huggingface.co/facebook/encodec_24khz)
DAC Kumar et al. (2024) 6 8 600 75 24kHz (https://github.com/descriptinc/descript-audio-codec/releases/download/0.0.4/weights_24khz.pth)
SpeechTokenizer Zhang et al. (2024b) 4 8 400 50 16kHz (https://huggingface.co/fnlp/AnyGPT-speech-modules/tree/main/speechtokenizer)
SQ-Codec Yang et al. (2024) 8 1 50 50 16kHz (https://huggingface.co/Dongchao/UniAudio/resolve/main/16k_50dim_9.zip)
SoundStream (ESPnet) Shi et al. (2024) 4 8 400 50 16kHz (https://huggingface.co/espnet/owsmdata_soundstream_16k_200epoch)
Speech-DAC (ESPnet) Shi et al. (2024) 4 8 400 50 16kHz (https://huggingface.co/ftshijt/espnet_codec_dac_large_v1.4_360epoch)
Mimi Défossez et al. (2024b) 4 32 100 12.5 16kHz (https://huggingface.co/kyutai/mimi)
BigCodec Xin et al. (2024) 1 1 80 80 16kHz (https://huggingface.co/Alethia/BigCodec/resolve/main/bigcodec.pt)
WavTokenizer Ji et al. (2024) 1 1 75 75 16kHz (https://huggingface.co/novateur/WavTokenizer-large-speech-75token)
TS3-Codec Wu et al. (2024a) 1 1 50 50 16kHz -
Table 4: VERSA demonstration on speech coding. Codecs with streaming capacities are marked with +++. The performance is evaluated on the LibriSpeech-test-clean set.
Codec kbps PESQ($\uparrow$) UTMOS($\uparrow$) DNSMOS (P.835)($\uparrow$) SPK-SIM($\uparrow$)
Encodec+ Défossez et al. (2024a) 6.00 2.77 3.09 2.96 0.72
DAC Kumar et al. (2024) 6.00 3.40 3.60 3.16 0.73
SpeechTokenizer Zhang et al. (2024b) 4.00 2.62 3.84 3.17 0.86
SQ-Codec Yang et al. (2024) 8.00 4.24 4.05 3.21 0.96
SoundStream (ESPnet) Shi et al. (2024) 4.00 2.86 3.61 3.13 0.89
Speech-DAC (ESPnet) Shi et al. (2024) 4.00 3.10 3.87 3.23 0.87
Mimi+ Défossez et al. (2024b) 4.40 3.38 3.92 3.18 0.85
BigCodec Xin et al. (2024) 1.04 2.68 4.11 3.26 0.62
WavTokenizer Ji et al. (2024) 0.98 1.88 3.77 3.18 0.60
TS3-Codec+ Wu et al. (2024a) 0.85 2.23 3.84 3.21 0.70
Ground Truth - - 4.09 3.18 -
Efficient Resource Management: Controlling external resources is a key consideration in the design of VERSA, particularly for scenarios involving the use of external resources such as pre-trained models, datasets, and reference files. To optimize performance and reduce redundant downloads, VERSA implements a robust cache control mechanism. This mechanism ensures that once a resource is downloaded, it is stored locally and can be reused across different metrics and evaluation runs. By centralizing resource management, VERSA minimizes network overhead and computational costs while maintaining consistency across evaluations.
Functional Interface: VERSA is designed with versatility and ease of use in mind, offering both command-line and Python function interfaces to accommodate diverse user preferences and workflows. While the command-line interface provides a straightforward way to execute evaluations, the Pythonic API enables seamless integration into custom scripts and larger pipelines, making it particularly useful for researchers and developers looking to automate or extend evaluations. To demonstrate this flexibility, we provide an interactive Google Colab notebook showcasing how VERSA can be effortlessly utilized within Python. [^5] https://colab.research.google.com/drive/11c0vZxbSa8invMSfqM999tI3MnyAVsOp This demonstration highlights the simplicity of configuring and executing evaluations programmatically, empowering users to leverage VERSA for both standard and highly customized use cases with minimal setup.
Community-driven Contribution: VERSA is designed not just as a toolkit, but as an evolving, community-driven platform that thrives on collaboration and shared innovation. To encourage and facilitate contributions, we provide clear and comprehensive guidelines for community members, available at https://github.com/wavlab-speech/versa/blob/main/contributing.md. These guidelines outline the process for proposing new features, adding metrics, reporting issues, and improving the existing codebase. By fostering an open and inclusive environment, VERSA invites researchers, developers, and practitioners from diverse backgrounds to contribute their expertise, ensuring that the toolkit remains up-to-date and relevant for a wide range of applications. This community-driven approach not only accelerates the development of VERSA but also ensures that it reflects the collective needs and insights of the global sound generation research community. Through active collaboration, VERSA aims to establish itself as a dynamic and enduring resource for the advancement of sound evaluation technologies.
Table 5: Detailed information of TTS models. AR and NAR stand for autoregressive and non-autoregressive architectures.
Model Feature Type Model Type Pre-trained Model Link
ESPnet-TTS Hayashi et al. (2021) Continuous NAR (https://huggingface.co/espnet/kan-bayashi_ljspeech_vits)
ESPnet-SpeechLM Discrete AR (https://huggingface.co/espnet/speechlm_tts_v1)
ChatTTS 2Noise (2024) Discrete AR (https://huggingface.co/2Noise/ChatTTS)
CosyVoice Du et al. (2024) Discrete AR (https://huggingface.co/model-scope/CosyVoice-300M)
EmotiVoice Youdao (2024) Continuous NAR (https://www.modelscope.cn/syq163/outputs.git)
MeloTTS Zhao et al. (2023) Continuous NAR (https://huggingface.co/myshell-ai/MeloTTS-English)
Parler-TTS Lacombe et al. (2024) Discrete AR (https://huggingface.co/parler-$t$s/parler-$t$s-mini-v1)
WhisperSpeech Collabora (2024) Discrete AR (https://huggingface.co/WhisperSpeech/WhisperSpeech/blob/main/t2s-v1.95-small-8lang.model)
VallE-X Zhang et al. (2023c) Discrete AR & NAR (https://huggingface.co/Plachta/VALL-E-X/resolve/main/vallex-checkpoint.pt)
Vall-E 2 Chen et al. (2024b) Discrete AR & NAR (https://huggingface.co/amphion/valle)
NaturalSpeech2 Shen et al. (2023) Continuous NAR (https://huggingface.co/amphion/naturalspeech2_libritts)
Table 6: VERSA demonstration on TTS. We use $\dagger$ for TTS systems with a fixed speaker.
Model UTMOS($\uparrow$) DNSMOS (P.835)($\uparrow$) SHEET-SSQA($\uparrow$) PLCMOS($\uparrow$) SPK-SIM($\uparrow$) TS-PESQ($\uparrow$) TS-SI-SNR($\uparrow$) TS-MOS($\uparrow$) Whisper-WER($\downarrow$)
ESPnet-TTS Hayashi et al. (2021) 4.06 3.05 4.02 4.51 - 3.61 26.55 4.25 5.01
ESPnet-SpeechLM 4.05 3.24 4.24 4.44 0.55 3.58 23.94 4.33 3.12
[…]
VallE-X Zhang et al. (2023c) 3.38 3.25 3.64 4.45 0.35 3.58 21.42 4.34 27.33
Vall-E 2 Chen et al. (2024b) 3.65 2.88 3.84 3.67 0.46 3.31 18.74 4.12 27.83
NaturalSpeech2 Shen et al. (2023) 2.35 2.24 3.16 4.27 0.32 2.27 10.85 3.62 4.62
Ground Truth 4.09 3.18 4.35 4.16 - 3.59 24.01 4.53 2.59
Appendix C Experimental Demonstration
This appendix includes the experimental results for the demonstration discussed in Section 4.
C.1 Codec Evaluation Demonstration
We demonstrate the usage of VERSA for speech coding evaluation with nine open-source models, including Encodec Défossez et al. (2024a), DAC Kumar et al. (2024), Mimi Défossez et al. (2024b), BigCodec Xin et al. (2024), SpeechTokenizer Zhang et al. (2024b), Wavtokenizer Ji et al. (2024), SQ-Codec Yang et al. (2024), TS3-Codec Wu et al. (2024a), and several ESPnet-Codec public checkpoints Shi et al. (2024). We use the test-clean set in Librispeech for evaluation Panayotov et al. (2015).
For all selected models, we utilize their public checkpoints along with the corresponding released inference scripts. Additionally, for multi-level codec models, we limit the codec levels to ensure alignment within a comparable bitrate range. Detailed inference hyperparameter configurations and the corresponding results are provided in Table 3.
Table 4 presents the evaluation results using the VERSA minimum installation. While detailed model discussions are beyond the scope of this paper, VERSA offers a versatile perspective for analyzing models and assessing their advantages. This capability significantly reduces the effort both model developers and end-users need to select codecs tailored to specific downstream applications. For example, the high pseudo-MOS scores (UTMOS and DNSMOS) observed in BigCodec indicate superior synthesis quality. However, the lower signal-level metrics (e.g., SDR and PESQ) suggest that the low bitrate leads to a substantial loss of signal detail.
While the current investigation serves as a demonstration with a small portion of the supported metrics in VERSA, we aim to conduct a more comprehensive analysis in future work, ensuring broader coverage across various aspects of sound generation.
Table 7: VERSA demonstration on speech enhancement. The performance is evaluated on the Voicebank-DEMAND test set.
Model UTMOS($\uparrow$) DNSMOS (P.835)($\uparrow$) SHEET-SSQA($\uparrow$) PESQ($\uparrow$) STOI($\uparrow$)
Noisy 3.11 2.52 3.57 1.97 0.92
MetricGAN+ Fu et al. (2021) 3.63 2.95 3.91 3.15 0.93
MTL-Mimic Bagchi et al. (2018) 3.86 3.03 4.06 3.01 0.95
SepFormer Subakan et al. (2021) 3.68 2.87 3.63 3.13 0.93
Table 8: Detailed pre-trained model links for the singing voice synthesis demonstration.
Table 9: VERSA demonstration on singing voice synthesis. The performance is evaluated on the Opencpop test set.
Model MCD($\downarrow$) F0-RMSE($\downarrow$) F0-CORR($\uparrow$) SHEET-SSQA($\uparrow$) SingMOS($\uparrow$) SPK-SIM($\uparrow$)
RNN Shi et al. (2021) 9.52 86.40 0.51 3.38 3.64 0.58
XiaoiceSing Lu et al. (2020) 9.54 69.05 0.57 3.13 3.53 0.57
DiffSinger Liu et al. (2022) 8.22 63.84 0.61 4.15 4.25 0.71
VISinger Zhang et al. (2022) 7.91 60.20 0.62 4.12 4.13 0.76
VISinger2 Zhang et al. (2023b) 7.89 59.53 0.62 4.19 4.21 0.62
TokSing Wu et al. (2024c) 8.10 61.59 0.62 3.73 3.86 0.67
VISinger2+ Yu et al. (2024) 8.73 55.26 0.66 4.32 4.22 0.68
Ground Truth - - - 4.49 4.79 -
Table 10: Detailed pre-trained model links for the music generation demonstration.
Table 11: VERSA demonstration on music generation.
Model FAD-LAION-CLAP($\downarrow$) KID-LAION-CLAP($\downarrow$) Density-LAION-CLAP($\uparrow$) Coverage-LAION-CLAP($\uparrow$)
AudioLDM2 Liu et al. (2023) 47.97 2.05 0.04 0.0058
MusicGen-large Copet et al. (2024) 20.06 0.49 0.07 0.0514
MusicLDM Chen et al. (2024a) 40.46 1.62 0.07 0.0118
Riffusion-v1Forsgren and Martiros (2022) 52.58 2.63 0.01 0.0024
StableAudioOpen Stability-AI (2023) 20.99 0.62 0.08 0.0674
C.2 TTS Evaluation Demonstration
Similar to the codec demonstration, we utilize a range of open-source TTS models for the TTS evaluation, including ESPnet-TTS Hayashi et al. (2021) [^6] We use the VITS model Kim et al. (2021) trained on LJSpeech Ito and Johnson (2017)., ESPnet-SpeechLM, [^7] https://github.com/espnet/espnet/tree/speechlm ChatTTS 2Noise (2024), CosyVoice Du et al. (2024), EmotiVoice Youdao (2024), MeloTTS Zhao et al. (2023), Parler-TTS Lacombe et al. (2024), WhisperSpeech Collabora (2024), VallE-X Zhang et al. (2023c), and Vall-E 2 Chen et al. (2024b). Specifically, we use the common LibriSpeech test-clean set for evaluation. For zero-shot TTS systems, we provide speaker prompts by randomly selecting a prompt segment from the target speaker. Table 5 provides a high-level categorization of feature types—discrete or continuous acoustic features—and model types, namely autoregressive or non-autoregressive. Inference is performed using the open-source pipelines provided by the model developers, with no additional hyperparameter tuning.
Some models, such as Parler-TTS and CosyVoice, struggled to interpret capitalized input text from LibriSpeech. To address this, we optionally applied a case-recovery preprocessing step for those models. Importantly, no post-filtering was applied to autoregressive models utilizing discrete tokens.
In Table 6, we present the performance of various TTS models. Since these models have been optimized using different frameworks and datasets, deriving definitive scientific conclusions is challenging. However, two potential implications can be drawn in this context:
  • There remains room for improvement in zero-shot TTS performance compared to fixed-speaker TTS systems.
  • Discrete token-based autoregressive modeling continues to be a popular approach. However, without sufficient post-filtering, it struggles to achieve stability comparable to that of non-autoregressive methods.
Building on this demonstration, we anticipate a more comprehensive analysis utilizing all 54 speech-related metrics in VERSA for existing TTS models in our future work. This analysis would further explore the current state of TTS development and address the challenges associated with existing evaluation metrics.
C.3 Speech Enhancement Evaluation Demonstration
We demonstrate the speech enhancement evaluation with three public available models, including SepFormer Subakan et al. (2021), MetricGAN Fu et al. (2021), and MTL-MIMIC Bagchi et al. (2018). We use VoiceBank-DEMAND as the test set Valentini-Botinhao (2017).
The evaluation of speech enhancement is demonstrated in Table 7. While this paper focuses on six specific metrics, we anticipate broader adoption of the task by leveraging the diverse set of 65 metrics available in VERSA.
C.4 Singing Synthesis Evaluation Demonstration
For the singing voice synthesis demonstration, we select a range of open-source singing voice synthesis models, including Naive-RNN Shi et al. (2021), XiaoiceSing Lu et al. (2020), VISinger Zhang et al. (2022), VISinger2 Zhang et al. (2023b), TokSing Wu et al. (2024c), VISinger2+ Yu et al. (2024), and DiffSinger Liu et al. (2022). [^8] The DiffSinger is trained with its own open-source repository Liu et al. (2022), where other models are trained with Muskit-ESPnet Shi et al. (2022); Wu et al. (2024b). The links for pre-trained models are listed in Table 8.
We demonstrate the usage of VERSA in singing voice synthesis in Table 9. Compared to speech quality assessment, singing voice-related analysis usually utilizes subjective evaluation due to the limited resources available. With the support of VERSA, we additionally support SHEET-SSQA Huang et al. (2024a) and SingMOS Tang et al. (2024), which consider singing voice synthesis, aligning a range of open-source singing synthesis models.
C.5 Music Generation Evaluation Demonstration
We utilize five models for music generation evaluation, including AudioLDM2 Liu et al. (2024), MusicGen Copet et al. (2024), MusicLDM Chen et al. (2024a), Riffusion-v1 Forsgren and Martiros (2022), StableAudioOpen Stability-AI (2023). Table 10 shows the pre-trained model links used for the demonstration. To generate music from each model, we utilized text prompts from MusicCaps, a dataset comprising expert-annotated captions of 30-second YouTube clips Agostinelli et al. (2023). We employed the rewritten, quality-neutral prompts provided in Gui et al. (2024). Additionally, ChatGPT was used to filter out prompts referring to non-instrumental music or those not typically associated with the text-to-music generation, such as tutorials or environmental sounds.
For evaluation, we utilize the final layer embedding of LAION-CLAP (Music pre-trained version) to compute the metrics Wu et al. (2023). [^9] https://github.com/LAION-AI/CLAP FMA-small dataset is used as the reference dataset for distributional analysis Defferrard et al. (2017).
Table 11 presents the evaluation results obtained using VERSA for music generation tasks. The results highlight the strengths and unique characteristics of different models. Specifically, MusicGen-large demonstrates robust capabilities in generating coherent and high-quality songs, showcasing its strength in musical structure and tonal consistency. On the other hand, StableAudioOpen excels in diversity, producing a wide range of musical styles and compositions. These results underscore the utility of VERSA in providing a detailed, multi-faceted evaluation of music generative models, enabling researchers to identify and compare model-specific strengths.